* A quality characteristic has a design specification (in cm) of 0.200 ± 0.020. If the actual process value of the quality characteristic is at the boundary of the tolerance on either side, the product will require a repair of $30 Find the value of k, and state the Taguchi loss function. What is the economic loss associated with x = 0.205?
media%2F2bb%2F2bbf969d-3ba5-41f6-bf72-f8
Answer #1

The Taguchi loss function is:-

L(x)=k(x-T)2

30=k(0.02)2

K=75000

The Taguchi loss function is:-

L(x)=75000(x-0.2)2

The loss associated with x=0.205

L(0.205)=75000(0.205-0.2)2

L=1.875